OpenStudy (confusedhampster):

How do I simplify this: (xy^3/x^-4y^4)^-2

OpenStudy (zenmo):

Try putting that problem on https://www.symbolab.com/

OpenStudy (zenmo):

it will show you "steps by steps" process on how to simplify it.

OpenStudy (mrnood):

or better still TRY solving here yourself with help

OpenStudy (zenmo):

just an option, while waiting for help.

OpenStudy (confusedhampster):

ok

OpenStudy (confusedhampster):

What do I do if there is a positive and negative exponent together?

OpenStudy (mrnood):

x/x^-4 * y^3/-4y^4 can you simplify each of those terms?

OpenStudy (confusedhampster):

ok, I'll try the first one.

OpenStudy (anonymous):

or you can go here http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

and it = 1 x10y14

OpenStudy (confusedhampster):

x^-4=1/x^4= x^4?

OpenStudy (confusedhampster):

I need it done step by step